OpenStudy (anonymous):

Three questions I'm confused on: P.S THE X2 ARE SQUARED! 1. Let f(x) = x + 7 and g(x) = x - 4. Find f(x) · g(x). x2 - 3x - 28 x2 - 3x - 11 x2 + 3x - 28 x2 + 3x - 11 2.Let f(x) = x2 + 3x - 4 and g(x) = x + 5. Find f(x) · g(x). x3 + 3x2 + 16x - 20 x3 + 5x2 + 14x - 20 x3 + 8x2 + 11x - 20 x3 + 9x2 + 19x - 20

OpenStudy (anonymous):

3. Let f(x) = 2x2 - 2x + 7 and g(x) = x2 + 9x - 6. Find f(x) - g(x). x2 - 11x + 13 x2 - 11x + 1 x2 + 7x + 1 x2 + 7x + 13

OpenStudy (anonymous):

For the first one I think its either c or d

OpenStudy (ankit042):

lol Maths is not luck based blind shooting. Try solving I will correct you if you make a mistake

OpenStudy (anonymous):

Okay. X times x = X2+ 7-4=3x..Then im lost from there

OpenStudy (anonymous):

Wait no, then 7 times 4=28?

OpenStudy (anonymous):

x2 + 3x - 28

OpenStudy (ankit042):

Yes see you know the stuff!

OpenStudy (anonymous):

:) okay, I'll try the next one

OpenStudy (anonymous):

so x2+x=x3 then 3x+5x=8x^2 then I know 4 times 5 =20, but I don't know how to get the middle number but i'm pretty sure its x3 + 8x2 + 11x - 20

OpenStudy (ankit042):

x2+x is NOT x3 that is true when you multiply x2*x=x3

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

so is x3 + 8x2 + 11x - 20 right?

OpenStudy (ankit042):

NO

OpenStudy (ankit042):

sry I was looking at the third question

OpenStudy (anonymous):

lol that's okay...so were did I mess up on number 2?

OpenStudy (anonymous):

I actually looked at it again, and know I think its x3 + 3x2 + 16x - 20

OpenStudy (ankit042):

( x2 + 3x - 4)(x + 5) = x2(x+5) +3x(x+5)-4(x+5) now expand and take all like terns together

OpenStudy (anonymous):

I have know idea how to do this.

OpenStudy (ankit042):

actually what you said before was correct! go with 8x^2 solution.

OpenStudy (anonymous):

OH OKAY :), and the last one I think is x2 - 11x + 1

OpenStudy (ankit042):

NO it is A +7-(-6)=+13

OpenStudy (anonymous):

So x2 + 7x + 13

OpenStudy (ankit042):

option A

OpenStudy (anonymous):

oh wait you meant x2 - 11x + 13 Okay! Thanks for all your help!

OpenStudy (ankit042):

np :)